# ECMSim: A high-performance web simulation of cardiac ECM remodeling through integrated ODE-based signaling and diffusion

### v1

@@ Abstract
Extracellular matrix (ECM) remodeling is central to a wide variety of healthy and diseased tissue processes. Unfortunately, predicting ECM remodeling under various chemical and mechanical conditions has proven to be excessively challenging, due in part to its complex regulation by intracellular and extracellular molecular reaction networks that are spatially and temporally dynamic. We introduce ECMSim, which is a highly interactive, real-time and web application designed to simulate heterogeneous matrix remodeling. The current model simulates cardiac scar tissue with configurable input conditions using a large-scale model of the cardiac fibroblast signaling network. Cardiac fibrosis is a major component of many forms of heart failure. ECMSim simulates over 1.3 million equations simultaneously in real time that include more than 125 species and more than 200 edges in each cell in a 100×100 spatial array (10,000 cells), which accounts for inputs, receptors, intracellular signaling cascades, ECM production, and feedback loops, as well as molecular diffusion. The algorithm is represented by a set of ordinary differential equations (ODEs) that are coupled with ECM molecular diffusion. The equations are solved on demand using compiled C++ and the WebAssembly standard. The platform includes brush-style cell selection to target a subset of cells with adjustable input molecule concentrations, parameter sliders to adjust parameters on demand, and multiple coupled real-time visualizations of network dynamics at multiple scales. Implementing ECMSim in standard web technologies enables a fully functional application that combines real-time simulation, visual interaction, and model editing. The software enables the investigation of pathological or experimental conditions, hypothetical scenarios, matrix remodeling, or the testing of the effects of an experimental drug(s) with a target receptor.
@@ Introduction
Extracellular matrix (ECM) remodeling represents one of the most computationally intensive and mechanistically complex biological processes in cardiovascular pathophysiology that require simultaneous integration of multi-scale reaction-diffusion dynamics, mechanochemical dynamics, and spatially heterogeneous cellular responses that span orders of magnitude in both temporal and spatial domains. 1, 2 The inherent difficulty of simulating real-time matrix remodeling is due to the complex interdependencies between collagen synthesis and degradation kinetics, the regulation of matrix metalloproteinases (MMPs), and mechanical feedback loops that arise from the changing stiffness of the tissue and the orientation of the fiber. These interdependencies have been explored in various studies that investigate the roles of MMPs and their regulatory mechanisms in ECM homeostasis, particularly in pathological conditions such as myocardial infarction (MI) 3 and idiopathic pulmonary fibrosis. 4 Moreover, the computational challenges posed by these dynamic interactions are well-documented, underscoring the need for advanced simulation techniques that can accommodate the intricacies of such biological systems. 5, 4 The clinical importance of dysregulation of ECM is highlighted by the fact that MI remains a of the leading causes of death worldwide, accounting for an estimated 16 million deaths annually, based on data from the World Health Organization (WHO). 6 Heart failure, which affects more than 64 million people worldwide, is characterized by significant matrix remodeling dynamics, driven in part by aging populations and improved survival rates from acute coronary events. 7, 8, 9 The association between ECM alterations and disease progression reinforces the urgency of addressing these remodeling challenges. 10
@@
Cardiac fibrosis plays an important role in heart failure, hypertrophic cardiomyopathy, and other heart illnesses. An excess of ECM proteins, especially collagens, is what makes this process happen. This changes the normal shape and function of the myocardium. 11, 12, 13 Cardiac fibroblasts are the primary cells that make the ECM, and they are responsive to multiple biophysical activities, like mechanical stress, inflammatory cytokines, growth factors, and neurohormonal signals. These signals work together in complicated ways to keep the balance between synthesis and degradation of the ECM. 14, 15 The complexity of fibroblast signaling pathways presents key challenges for scientific studies, especially in matrix remodeling with multi-species diffusion. These pathways are interconnected and include transforming growth factor- (TGF-) signaling, mitogen-activated protein kinase (MAPK) cascades, mechanotransduction pathways, and inflammatory mediator responses. 15, 16 The dynamic interactions among these signaling pathways occur over various timescales, from immediate receptor activation to prolonged transcriptional changes that can last several hours to days. Furthermore, the spatial heterogeneity in cellular responses makes it more complex with the complexity of these signaling networks, as does the intercellular communication facilitated by paracrine signaling. 17 Traditional experimental methods often fall short in effectively capturing these multifaceted interactions in especially in extracellular simulation with multiple-molecular diffusion with feedback loops.
@@
In response of these shortcomings, computational modeling has become an important tool in cardiac fibrosis studies because it lets researchers combine different experimental results into complete models. These kinds of mathematical systems not only help come up with new hypotheses, but they also provide a way to test possible treatments. 18, 19 Nonetheless, numerous current computational models of cardiac fibroblast signaling exhibit considerable deficiencies. A prevalent deficiency is that many of these models don’t have the spatial resolution needed to accurately model how cells interact with each other and how tissues are organized. Most computational tools demand special software and knowledge, which makes it harder for more researchers to use them. 12, 20 Recent improvements in web technologies, including WebAssembly, show promise for solving these problems. This technology lets web browsers do high-performance calculations at speeds that are similar to those of native apps, while also being easy to use and working on all platforms. 21 These features are especially useful for scientific computing programs that need a lot of numerical analysis, which is very important in an investigation of cardiac fibrosis. This paper introduces ECMSim, a new web-based platform that simulates the signaling network of cardiac fibroblasts. This platform gets around the problems with current tools by using a full and system-level model of fibroblast signaling in an easy-to-use interface. 22 ECMSim consists of ordinary differential equations (ODEs) for mathematical representations of all the species organized into various functional modules that span receptor activation, inhibition, intracellular signaling cascades, transcriptional regulation, and intercellular feedback mechanisms. The combination of all the ODEs in all the cells are distributed in the grid with different altering input molecules and molecular diffusion in the spatial matrix. The computational engine harnesses optimized C++ code compiled to WebAssembly, delivering real-time performance for simulations involving extensive cellular arrangements. 18
@@
By establishing a robust and user-friendly platform for real-time simulating cardiac fibroblast interactions with temporal and spatial dyanmics ECMSim significantly contributes to advancing the understanding of the molecular mechanisms underlying cardiac fibrosis and facilitates the exploration of novel therapeutic strategies. ECMSim is not limited to cardiac fibrotic matrix remodeling. It provides the foundation to model the advanced reaction-diffusion dynamics of specific applications, including cancer micro-environment, cell cross-talk, regenerative biology, wound healing, and drug discovery.
@@
[width=1]Fig-1.pdf Network architecture of the cardiac fibroblast signaling model implemented in ECMSim. The network integrates ten input molecules (E2, IL6, NP, TGF-, IL1, TNF-, NE, PDGF, ET1, AngII, and tension) that activate membrane receptors and trigger intracellular signaling cascades. Key signaling modules include MAPK pathways (ERK, p38, JNK), PI3K-Akt-mTORC signaling, Rho GTPase mechanotransduction, and calcium-dependent pathways. These converge on transcription factors (STAT, Smad3, CREB, AP1, NFAT, SRF, YAP, NF-B) that regulate expression of ECM molecules including collagens (proCI, proCIII), matrix metalloproteinases (proMMP1, 2, 3, 8, 9, 12, 14), tissue inhibitors (TIMP1, TIMP2), matricellular proteins (CTGF, periostin, fibronectin, tenascin-C, thrombospondin-4, osteopontin), and other ECM regulators (SMA, PAI1, EDAFN). Feedback molecules (TGF-fb, AngIIfb, IL6fb, ET1fb, tensionfb) enable autocrine and paracrine signaling. Arrows indicate activation, bar-headed lines indicate inhibition, and “AND” gates represent cooperative regulatory interactions requiring multiple inputs for activation.
@@
The ECMSim platform models cardiac fibroblast signaling through a comprehensive system of ordinary differential equations (ODEs) describing the temporal evolution of molecular species concentrations across a spatially-resolved tissue grid. The complete system encompasses all the molecular species per cell distributed across a cellular grid, yielding intracellular variables. Including four diffusible feedback molecules (TGF-fb, AngIIfb, IL6fb, and ET1fb) across the spatial domain adds additional variables, resulting in coupled differential equations solved simultaneously at each time step.
@@
ensuring concentrations remain within the normalized range representing physiologically relevant activity levels.
[width=1]Fig-2.pdf Representation of the integrated cardiac fibroblast signaling model implemented in ECMSim, illustrating the key components and their interactions. The model encompasses; (A) Single feedback loop: representation of how feedback loops connected in ODE system; (B) ODE-based signaling pathway: crosstalk between signaling modules including MAPK cascades (ERK, p38, JNK), PI3K-Akt-mTOR, Rho GTPase, and mechanotransduction pathways; (C) Matrix remodeling: include cell-cell crosstalk with feedback molecules and ECM molecule diffusion; (D) Diffusion dynamics: spatial diffusion to neighboring cells, demonstrating the multi-scale integration from molecular interactions to tissue-level coordination according to the given conditions in the ECMSim platform.
@@
The signaling network (Figure 1) architecture based on previous mechnistic models that consist of molecular species arranged into functionally distinct yet interconnected modules that mirror biological organization while facilitating extensive crosstalk among pathways. 23, 13, 24 The input processing module handles ten primary signaling molecules including TGF-, angiotensin II (AngII), interleukin-6 (IL6), interleukin-1 (IL1), tumor necrosis factor- (TNF-), norepinephrine (NE), platelet-derived growth factor (PDGF), endothelin-1 (ET1), natriuretic peptides (NP), and estrogen (E2). Each input signal incorporates both external stimuli and intercellular feedback according to:
where represents user-defined input concentrations applied to selected cell populations through the brush selection interface, represents diffusible signals from neighboring cells, and rate constants , , and govern input processing, feedback sensitivity, and degradation, respectively (Figure 2A).
@@
[width=1]Fig-3.pdf Main areas and selective options of the first version of ECMSim application: Main heatmap shows concentration of the main selected molecule in the matrix. The application enables real-time modification of simulation parameters, allowing users to alter the primary molecular species and ODE parameters during active simulation execution. Dynamic selection of the main molecule automatically updates both spatial heatmap visualizations and temporal expression profiles, providing immediate feedback on parameter modifications without requiring simulation restart.
@@
ECMSim implements spatial organization (Figure 2) through a regular two-dimensional grid of individual cells, representing a tissue region of approximately 1 cm² when scaled to physiological dimensions. Each grid location contains a complete representation of the intracellular signaling network, creating a spatially-resolved model with over 1.3 million individual molecular species variables distributed across the computational domain.
Spatial coupling between cells is implemented through the diffusion of five key feedback molecules that mediate paracrine signaling and coordinate tissue-level responses. The spatial evolution follows a discrete diffusion equation using finite difference approximation on the cellular grid:
where represents the concentration of feedback molecule at grid position , denotes the eight-connected neighborhood of cell , is the diffusion coefficient, represents the cellular production rate dependent on the intracellular signaling state, and represents extracellular degradation or clearance.
@@
The five feedback molecules include secreted TGF- that acts in autocrine and paracrine modes to amplify fibrotic responses, angiotensin II produced through local renin-angiotensin system activation, IL6 that coordinates inflammatory responses, endothelin-1 that propagates vasoconstrictor and pro-fibrotic signals, and mechanical tension transmitted through cell-matrix connections enabling mechanochemical signal propagation.
Extracellular matrix molecule concentrations evolve according to production from activated fibroblasts and limited spatial diffusion reflecting their larger molecular size and tendency for local deposition:
where represents ECM molecule at position , is the corresponding intracellular precursor concentration, and ECM molecules diffuse at 20% the rate of feedback molecules to reflect their distinct biophysical properties.
@@
The time step is dynamically adjusted based on the maximum rate of change across all molecular species to ensure numerical stability. While higher-order integration methods such as Runge-Kutta algorithms offer improved accuracy, the forward Euler approach provides several advantages for interactive simulation applications, including computational efficiency and numerical stability for the specific rate constants employed, with predictable execution times that facilitate real-time visualization updates. All molecular concentrations are subject to physiological constraints at each integration step, ensuring values remain within normalized bounds, representing meaningful biological activity levels. The spatial derivatives in the diffusion equations are approximated using second-order finite difference methods with periodic boundary conditions to minimize edge effects during typical simulation scenarios. The model incorporates eight primary rate constants that can be adjusted (from the default set value) in real-time through the user interface: input processing rate (), feedback sensitivity (), degradation rate (), receptor activation rate (), inhibition rate (), activation rate (), ECM production rate (), and diffusion rate ().
@@
The computational core of ECMSim is implemented in C++ using object-oriented design principles with optimized data structures, including unordered maps for efficient molecular species lookup and vector arrays for spatial grid representation. Memory allocation is managed through pre-allocation strategies that minimize dynamic memory operations during simulation execution. The simulation state is organized to maximize cache locality by storing related molecular species in contiguous memory locations. The C++ simulation code is compiled to WebAssembly using the Emscripten toolchain with aggressive optimization flags including function inlining, loop unrolling, and dead code elimination. The resulting WebAssembly binary achieves computational performance within approximately 20% of equivalent native C++ code while maintaining complete cross-platform compatibility across modern web browsers. The WebAssembly linear memory model provides direct access to simulation data from JavaScript visualization code, eliminating costly data serialization and transfer operations. Memory management within the WebAssembly environment utilizes several optimization strategies: simulation state organization for cache locality, temporary variable reuse across integration steps, and leveraging the linear memory model for efficient data access. The platform executes efficiently on standard computing hardware, requiring only a modern web browser with WebAssembly support for typical simulations involving the full 1.36 million variable system.

### v2

# ECMSim: A high-performance interactive web application for real-time spatiotemporal simulation of cardiac ECM signaling and diffusion

@@ ABSTRACT
Extracellular matrix (ECM) remodeling is central to a wide variety of healthy and diseased tissue processes. Unfortunately, predicting ECM remodeling under various chemical and mechanical conditions has proven to be excessively challenging, due in part to its complex regulation by intracellular and extracellular molecular reaction networks that are spatially and temporally dynamic. We introduce ECMSim, which is a highly interactive, real-time, and web application designed to simulate heterogeneous matrix remodeling. The current model simulates cardiac scar tissue with configurable input conditions using a large-scale model of the cardiac fibroblast signaling network. Cardiac fibrosis is a major component of many forms of heart failure. ECMSim solves 1.37 million coupled ordinary differential equations (ODEs) and executes approximately 4.84 million operations per time step in real time, encompassing 137 molecular species and 259 regulatory interactions per cell across a 100100 spatial array (10,000 cells), which accounts for inputs, receptors, intracellular signaling cascades, ECM production, feedback loops, and molecular diffusion. The algorithm is represented by a set of ODEs that are coupled with ECM molecular diffusion. The equations are solved on demand using compiled C++ and the WebAssembly standard. The platform includes brush-style cell selection to target a subset of cells with adjustable input molecule concentrations, parameter sliders to adjust parameters on demand, and multiple coupled real-time visualizations of network dynamics at multiple scales. Implementing ECMSim in standard web technologies enables a fully functional application that combines real-time simulation, visual interaction, and model editing. The software enables the investigation of pathological or experimental conditions, hypothetical scenarios, matrix remodeling, or the testing of the effects of an experimental drug(s) with a target receptor.
@@ INTRODUCTION
Extracellular matrix (ECM) remodeling represents one of the most computationally intensive and mechanistically complex biological processes in cardiovascular pathophysiology that require simultaneous integration of multi-scale reaction-diffusion dynamics, mechanochemical dynamics, and spatially heterogeneous cellular responses that span orders of magnitude in both temporal and spatial domains. 41, 22 The inherent difficulty of simulating real-time matrix remodeling is due to the complex interdependencies between collagen synthesis and degradation kinetics, the regulation of matrix metalloproteinases (MMPs), and mechanical feedback loops that arise from the changing stiffness of the tissue and the orientation of the fiber. These interdependencies have been explored in various studies that investigate the roles of MMPs and their regulatory mechanisms in ECM homeostasis, particularly in pathological conditions such as myocardial infarction (MI) 4 and idiopathic pulmonary fibrosis. 30 Moreover, the computational challenges posed by these dynamic interactions are well-documented, underscoring the need for advanced simulation techniques that can accommodate the intricacies of such biological systems. 47, 30 The clinical importance of dysregulation of ECM is highlighted by the fact that MI remains one of the leading causes of death worldwide, accounting for an estimated 16 million deaths annually, based on data from the World Health Organization (WHO). 25 Heart failure, which affects more than 64 million people worldwide, is characterized by significant matrix remodeling dynamics, driven in part by aging populations and improved survival rates from acute coronary events. 5, 3, 20 The association between ECM alterations and disease progression reinforces the urgency of addressing these remodeling challenges. 7
@@
Cardiac fibrosis plays an important role in heart failure, hypertrophic cardiomyopathy, and other heart illnesses. An excess of ECM proteins, especially collagens, is what makes this process happen. This excessive ECM accumulation changes the normal shape and function of the myocardium. 27, 12, 33 Cardiac fibroblasts are the primary cells that make the ECM, and they are responsive to multiple biophysical activities, like mechanical stress, inflammatory cytokines, growth factors, and neurohormonal signals. These signals work together in complicated ways to keep the balance between synthesis and degradation of the ECM. 35, 36 The complexity of fibroblast signaling pathways presents key challenges for scientific studies, especially in matrix remodeling with multi-species diffusion. These pathways are interconnected and include transforming growth factor- (TGF-) signaling, mitogen-activated protein kinase (MAPK) cascades, mechanotransduction pathways, and inflammatory mediator responses. 36, 19 The dynamic interactions among these signaling pathways occur over various timescales, from immediate receptor activation to prolonged transcriptional changes that can last several hours to days. Furthermore, the spatial heterogeneity in cellular responses adds additional complexity, as does the intercellular communication facilitated by paracrine signaling. 21 Traditional experimental approaches are limited in their ability to capture the full complexity of multi-pathway signaling dynamics combined with spatial molecular diffusion and feedback regulation. Computational modeling provides a complementary approach that can integrate these multiple interacting processes within a unified quantitative framework.
@@
In response of these shortcomings, computational modeling has become an important tool in cardiac fibrosis studies because it lets researchers combine different experimental results into complete models. These kinds of mathematical systems not only help come up with new hypotheses, but they also provide a way to test possible treatments. 42, 17 Nonetheless, numerous current computational models of cardiac fibroblast signaling exhibit considerable deficiencies. A prevalent deficiency is that many of these models don’t have the spatial resolution needed to accurately model how cells interact with each other and how tissues are organized. Most computational tools demand special software and knowledge, which makes it harder for more researchers to use them. 12, 2 Recent improvements in web technologies, including WebAssembly, show promise for solving these problems. This technology lets web browsers do high-performance calculations at speeds that are similar to those of native apps, while also being easy to use and working on all platforms. 44 These features are especially useful for scientific computing programs that need a lot of numerical analysis, which is very important in the investigation of cardiac fibrosis. This paper introduces ECMSim, a new web-based platform that simulates the signaling network of cardiac fibroblasts. This platform addresses the limitations of existing tools, including the lack of spatial resolution for cell-cell communication, the absence of real-time interactivity for parameter exploration, and the requirement for specialized software installation that limits accessibility, by using a comprehensive, system-level model that integrates all major signaling pathways implicated in cardiac fibroblast activation, from receptor-level inputs through intracellular signaling cascades to ECM protein outputs, within a single, unified computational framework in an easy-to-use interface. 15 ECMSim consists of ordinary differential equations (ODEs) for mathematical representations of all the species organized into various functional modules that span receptor activation, inhibition, intracellular signaling cascades, transcriptional regulation, and intercellular feedback mechanisms. The combination of all the ODEs in all the cells are distributed in the grid with different altering input molecules and molecular diffusion in the spatial matrix. The computational engine harnesses optimized C++ code compiled to WebAssembly, delivering real-time performance for simulations involving large-scale spatial tissue grids containing 10,000 individual cells, each with a complete intracellular signaling network. 42
ECMSim’s signaling network was informed by the logic-based ODE model of cardiac fibroblast signaling developed by Zeigler et al., 45, 46 which demonstrated that a single-cell model could predict context-dependent drivers of myofibroblast differentiation with approximately 80% accuracy across 82 input-output relationships. This mechanistic model has been well validated and extended in subsequent studies, including predictions of mechano-adaptive infarct therapies, 34 computational screens for sex-specific drug effects in cardiac fibroblasts, 43 and hierarchical molecular language models for signaling network analysis. 14 Related models by Rogers et al. 33 and Rikard et al. 32 have explored fibroblast mechanotransduction and ECM regulation, respectively. However, these models are limited to single-cell representations without spatial coupling, lack real-time interactivity, and require specialized software for execution. ECMSim extends this foundation in several key ways: (1) expansion from a single-cell model to a spatially-resolved tissue grid with reaction-diffusion coupling; (2) implementation of real-time WebAssembly-based computation enabling interactive browser-based simulation; and (3) feedback loops enabling autocrine and paracrine signaling between cells.
@@
By establishing a robust and user-friendly platform for real-time simulating cardiac fibroblast interactions with temporal and spatial dynamics, ECMSim significantly contributes to advancing the understanding of the molecular mechanisms underlying cardiac fibrosis and facilitates the exploration of novel therapeutic strategies. ECMSim is not limited to cardiac fibrotic matrix remodeling. It provides the foundation to model the advanced reaction-diffusion dynamics of specific applications, including cancer micro-environment, cell cross-talk, regenerative biology, wound healing, and drug discovery.
@@
Network architecture of the cardiac fibroblast signaling model implemented in ECMSim. The network integrates ten input molecules (E2, IL6, NP, TGF-, IL1, TNF-, NE, PDGF, ET1, AngII, and tension) that activate membrane receptors and trigger intracellular signaling cascades. Key signaling modules include MAPK pathways (ERK, p38, JNK), PI3K-Akt-mTORC signaling, Rho GTPase mechanotransduction, and calcium-dependent pathways. These converge on transcription factors (STAT, Smad3, CREB, AP1, NFAT, SRF, YAP, NF-B) that regulate expression of ECM molecules including collagens (proCI, proCIII), matrix metalloproteinases (proMMP1, 2, 3, 8, 9, 12, 14), tissue inhibitors (TIMP1, TIMP2), matricellular proteins (CTGF, periostin, fibronectin, tenascin-C, thrombospondin-4, osteopontin), and other ECM regulators (-SMA, PAI1, EDAFN). Feedback molecules (TGF-, AngII, IL6, ET1, tension) enable autocrine and paracrine signaling. Arrows indicate activation, bar-headed lines indicate inhibition, and “AND” gates represent cooperative regulatory interactions requiring multiple inputs for activation.
@@
The ECMSim platform models cardiac fibroblast signaling through a comprehensive system of ordinary differential equations (ODEs) describing the temporal evolution of molecular species concentrations across a spatially-resolved tissue grid. The complete system encompasses 132 molecular species per cell (115 intracellular signaling species and 17 extracellular ECM species) distributed across a cellular grid, yielding intracellular variables. Including five diffusible feedback signals (TGF-, AngII, IL6, ET1, and tension) across the spatial domain adds additional variables, resulting in coupled differential equations solved simultaneously at each time step. The signaling network contains 259 regulatory interactions (edges), of which approximately 216 are activating and 43 are inhibitory. Table 1 summarizes the distribution of molecular species across functional modules.
| Module | Species | Examples |
| Ligands/Inputs | 11 | AngII, TGF-, IL6, ET1, tension |
| Receptors | 10 | AT1R, TGFB1R, ETAR, PDGFR, gp130 |
| Second messengers | 7 | ROS, DAG, cAMP, cGMP, Ca2+ |
| Kinases/Phosphatases | 5 | PKA, PKC, PKG, calcineurin, PP1 |
| MAPK pathway | 11 | Ras, Raf, MEK1, ERK, p38, JNK |
| PI3K-Akt-mTOR | 6 | PI3K, Akt, mTORC1, mTORC2, p70S6K |
| Rho/ROCK & cytoskeleton | 20 | Rho, ROCK, FAK, Src, YAP, Factin |
| Transcription factors | 8 | NF-B, AP1, STAT, Smad3, SRF, NFAT |
| Estrogen signaling | 5 | ER, ER, GPR30, CyclinB1 |
| Additional regulatory | 14 | AGT, ACE, BAMBI, -SMA, LOX |
| ECM precursors (intracellular) | 17 | proCI, proCIII, proMMP2, TIMP1 |
| Intracellular subtotal | 115 | |
| ECM molecules (extracellular) | 17 | proCI, fibronectin, proMMP9, TIMP2 |
| Feedback molecules | 5 | TGF-, AngII, tension |
| Total per cell | 137 | |
| Total on grid () |
@@
ensuring concentrations remain within the normalized range , where 0 represents an inactive state and 1 represents maximal physiological activity. This clamping serves as a numerical safeguard against transient overshoots or negative values that the forward Euler integration can produce during a single time step, particularly when multiple strong activation signals converge simultaneously, while also providing a consistent and intuitive framework for comparing activity levels across diverse molecular species with vastly different absolute concentrations.
(A) Single feedback loop: TGF- input activates its receptor (TGF-R), which triggers intracellular signaling cascades that produce both ECM molecules and the feedback molecule TGF-; the feedback molecule diffuses to neighboring cells, completing the autocrine/paracrine loop. (B) ODE-based signaling model: hierarchical network architecture showing crosstalk between signaling modules including MAPK cascades (ERK, p38, JNK), PI3K-Akt-mTOR, Rho GTPase, and mechanotransduction pathways, from input receptors (top) through intracellular kinases to ECM outputs (bottom). (C) Matrix remodeling: the cellular grid representation with cell selection (cells 1 and 2 for temporal tracking), brush-based spatial selection of cell subpopulations, and bidirectional ECM molecule exchange between neighboring cells. (D) Diffusion dynamics: four representative simulation snapshots showing the temporal evolution of molecular concentrations from random initialization (top-left, iteration 0) through intermediate diffusion (top-right, iteration 300 without brush-selected input) to the equilibrated state without stimulation (bottom-left, iteration 150) and with localized stimulation (bottom-right, iteration 300), demonstrating how brush-applied inputs drive spatially heterogeneous concentration patterns. (E) Vector field streamlines illustrating the direction and magnitude of diffusive flux for four discrete concentration sources (points a, b, c, d) placed at different grid locations. (F) Gradient field showing the resulting two-dimensional concentration distribution from the four sources, with the color scale indicating normalized concentration (0–1). (G) Three-dimensional surface visualization of the same concentration field, where peak height corresponds to local concentration at each source location, illustrating how multiple diffusion sources interact to form the composite spatial concentration landscape.
@@
The signaling network (Fig. 1) architecture organizes molecular species into functionally distinct yet interconnected modules (receptors, second messengers, kinases, transcription factors, and ECM outputs) that reflect the hierarchical organization of biological signaling pathways, from extracellular inputs through intracellular processing to effector outputs, while facilitating extensive crosstalk among pathways. The input processing module handles ten primary signaling molecules including TGF-, angiotensin II (AngII), interleukin-6 (IL6), interleukin-1 (IL1), tumor necrosis factor- (TNF-), norepinephrine (NE), platelet-derived growth factor (PDGF), endothelin-1 (ET1), natriuretic peptides (NP), and estrogen (E2). Each input signal incorporates both external stimuli and intercellular feedback according to:
where represents user-defined input concentrations applied to selected cell populations through the brush selection interface, represents diffusible signals from neighboring cells, and rate constants , , and govern input processing, feedback sensitivity, and degradation, respectively (Fig. 2A).
@@
The interface comprises several interconnected panels. Left: Molecule selection sidebar listing all 17 ECM molecules and 4 feedback molecules available for spatial visualization. Top: The “Select molecule” dropdown determines which molecular species is displayed on the main heatmap; the input molecule concentration sliders (AngII, TGF-, IL6, tension, IL1, TNF-, NE, PDGF, ET1, NP, E2) set stimulus levels applied exclusively to brush-selected cells. Center: The brush selection tool enables spatial targeting of cell subpopulations with adjustable brush size (1–15); simulation controls (Start, Stop, Step, Reset) govern execution; the main heatmap displays real-time concentration dynamics with iteration count, simulation time, and concentration range. Export buttons allow saving the heatmap as SVG or PNG. Bottom: The cell tracking tool enables selection of up to 8 individual cells for temporal monitoring, with their concentration trajectories plotted in real time; tracked cell plots can also be exported. Bottom-left: The ODE parameters panel provides real-time adjustment of all nine primary rate constants (time step, input, feedback, degradation, receptor, inhibition, activation, production, and diffusion rates) via interactive sliders.
@@
ECMSim implements spatial organization (Fig. 2) through a regular two-dimensional grid of individual cells, where each grid site represents a single cardiac fibroblast. The grid spacing m corresponds to the typical diameter of a cardiac fibroblast (10–20 m), so the full grid represents a tissue region of approximately 1 mm 1 mm. Each grid location contains a complete representation of the intracellular signaling network, creating a spatially-resolved model with over 1.3 million individual molecular species variables distributed across the computational domain. For larger tissue regions, the grid dimensions can be increased (e.g., or ), with computational cost scaling linearly with the number of grid cells.
Spatial coupling between cells is implemented through the diffusion of five key feedback signaling molecules that mediate paracrine signaling and coordinate tissue-level responses. These are small, rapidly diffusing molecules, each with its own diffusion coefficient , reflecting their distinct molecular sizes and tissue transport properties. The spatial evolution follows a discrete diffusion equation using finite difference approximation on the cellular grid:
where represents the concentration of feedback molecule at grid position , denotes the eight-connected neighborhood of cell , is the molecule-specific diffusion coefficient, represents the cellular production rate that depends on the intracellular signaling state of the cell at position (i.e., is a function of relevant intracellular precursor concentrations), and represents extracellular degradation or clearance.
@@
The five feedback molecules include secreted TGF- that acts in autocrine and paracrine modes to amplify fibrotic responses, angiotensin II produced through local renin-angiotensin system activation, IL6 that coordinates inflammatory responses, endothelin-1 that propagates vasoconstrictor and pro-fibrotic signals, and biomechanical signal (tension) an indirect way to capture force transmission that can mechanically stimulate neighboring cells across a distance. In contrast to the feedback signaling molecules described above, ECM and ECM-binding molecules (collagens, fibronectin, periostin, etc.) are large, locally deposited proteins. Their concentrations evolve according to production from activated fibroblasts and limited spatial diffusion reflecting their larger molecular size and tendency for local deposition:
where represents ECM molecule at position , is the corresponding intracellular precursor concentration, and ECM molecules diffuse at 20% the rate of feedback molecules to reflect their distinct biophysical properties. The 0.2 scaling factor is grounded in biophysical evidence: ECM structural proteins such as collagens (100–400 kDa) and fibronectin (440 kDa) are substantially larger than feedback signaling molecules such as TGF- (25 kDa), IL6 (21 kDa), and AngII (1 kDa). According to Milo and Phillips, 29 diffusion coefficients decrease by roughly one order of magnitude per order-of-magnitude increase in molecular weight, yielding a ratio of approximately 0.1–0.2 for large versus small proteins. This size-dependent hindrance has been directly measured in collagen gels using fluorescence recovery after photobleaching (FRAP) 31 and modeled through steric, hydrodynamic, and electrostatic interactions in the ECM. 39
@@
The time step is dynamically adjusted based on the maximum rate of change across all molecular species to ensure numerical stability. While higher-order integration methods such as Runge-Kutta algorithms offer improved accuracy, the forward Euler approach provides several advantages for interactive simulation applications, including computational efficiency and numerical stability for the specific rate constants employed, with predictable execution times that facilitate real-time visualization updates. All molecular concentrations are subject to physiological constraints at each integration step, ensuring values remain within normalized bounds, representing meaningful biological activity levels. The spatial derivatives in the diffusion equations are approximated using second-order finite difference methods with periodic boundary conditions to minimize edge effects during typical simulation scenarios.
Parameter estimation and model calibration
The model incorporates eight primary rate constants that can be adjusted (from the default set value) in real-time through the user interface, as summarized in Table 2. As the model uses normalized concentrations in the range representing fractional activation levels, all rate constants have units of inverse time (time step-1). In the discrete diffusion formulation (Eq. 6), the grid spacing is already absorbed into the diffusion coefficient through the finite difference discretization, so .
| Parameter | Description | Default | Units |
| Input processing rate | 1.0 | time step-1 | |
| Feedback sensitivity | 0.5 | time step-1 | |
| Degradation rate | 0.1 | time step-1 | |
| Receptor activation rate | 2.0 | time step-1 | |
| Inhibition rate | 0.5 | time step-1 | |
| Activation rate | 1.0 | time step-1 | |
| ECM production rate | 0.01 | time step-1 | |
| Diffusion rate | 0.25 | time step-1 |
@@
The computational core of ECMSim is implemented in C++ using object-oriented design principles with optimized data structures, including unordered maps for efficient molecular species lookup and vector arrays for spatial grid representation. Memory allocation is managed through pre-allocation strategies that minimize dynamic memory operations during simulation execution. The simulation state is organized to maximize cache locality by storing related molecular species in contiguous memory locations. The C++ simulation code is compiled to WebAssembly using the Emscripten toolchain with aggressive optimization flags including function inlining, loop unrolling, and dead code elimination. The resulting WebAssembly binary achieves computational performance within approximately 20% of equivalent native C++ code while maintaining complete cross-platform compatibility across modern web browsers. The WebAssembly linear memory model provides direct access to simulation data from JavaScript visualization code, eliminating costly data serialization and transfer operations. Memory management within the WebAssembly environment utilizes several optimization strategies: simulation state organization for cache locality, temporary variable reuse across integration steps, and leveraging the linear memory model for efficient data access. The platform executes efficiently on standard computing hardware, requiring only a modern web browser with WebAssembly support for typical simulations involving the full 1.37 million variable system.
Performance benchmarks
Performance profiling of ECMSim was conducted using Chrome DevTools on a MacBook Pro (14-inch, 2021; Apple M1 Max; 32 GB RAM; Google Chrome). Table 3 summarizes the key performance metrics measured during real-time simulation of the full grid with 1.37 million coupled ODEs.
| Metric | Measured Value |
| Computational scale | |
| Total coupled ODEs | |
| Spatial grid | (10,000 cells) |
| Species per cell | 137 |
| Regulatory interactions (edges) per cell | 259 |
| Operations per time step | |
| ODE rate evaluations () | 1,370,000 |
| Euler integration updates | 1,490,000 |
| Feedback diffusion ( neighbors) | 450,000 |
| ECM diffusion ( neighbors) | 1,530,000 |
| Total operations per time step | |
| Memory and runtime | |
| Total memory footprint (browser tab) | 523 MB |
| JS Heap memory | 37.2–42.2 MB |
| CPU time: Scripting (ODE solver) | 99.4% |
| CPU time: Painting (canvas rendering) | 0.25% |
| CPU time: Rendering (layout) | 0.17% |
| CPU time: System overhead | 0.15% |
| CPU utilization | 105% (single-core) |
| Minimum recommended RAM | 4 GB |
| Simulation mode | Real-time interactive |
| Browser compatibility | Chrome, Firefox, Safari |
Each simulation time step executes approximately 4.84 million operations, comprising 1.37 million ODE rate evaluations, 1.49 million Euler integration updates, and 1.98 million diffusion neighbor computations across the feedback (5 molecules) and ECM (17 molecules) spatial grids. The total browser memory footprint of approximately 523 MB (including WebAssembly linear memory, JavaScript heap, and canvas rendering buffers) is modest for a system of this scale, and the JavaScript heap alone requires only 37–42 MB. The CPU profiling confirms that over 99% of computation time is dedicated to the ODE integration, with negligible overhead for visualization rendering (0.5%), indicating efficient allocation of computational resources to the numerical simulation core. The platform has been tested across Chrome, Firefox, and Safari, with Chrome providing the best WebAssembly performance due to its optimized V8 engine.